### v1

# Wasserstein Convergence of Score-based Generative Models under Semiconvexity and Discontinuous Gradients

Abstract
Score-based Generative Models (SGMs) approximate a data distribution by perturbing it with Gaussian noise and subsequently denoising it via a learned reverse diffusion process. These models excel at modeling complex data distributions and generating diverse samples, achieving state-of-the-art performance across domains such as computer vision, audio generation, reinforcement learning, and computational biology. Despite their empirical success, existing Wasserstein-2 convergence analysis typically assume strong regularity conditions—such as smoothness or strict log-concavity of the data distribution—that are rarely satisfied in practice.
In this work, we establish the first non-asymptotic Wasserstein-2 convergence guarantees for SGMs targeting semiconvex distributions with potentially discontinuous gradients. Our upper bounds are explicit and sharp in key parameters, achieving optimal dependence of $O(\sqrt{d})$ on the data dimension $d$ and convergence rate of order one. The framework accommodates a wide class of practically relevant distributions, including symmetric modified half-normal distributions, Gaussian mixtures, double-well potentials, and elastic net potentials. By leveraging semiconvexity without requiring smoothness assumptions on the potential such as differentiability, our results substantially broaden the theoretical foundations of SGMs, bridging the gap between empirical success and rigorous guarantees in non-smooth, complex data regimes.
1 Introduction
Score-based Generative Models (SGMs), also known as diffusion-based generative models [55, 57, 54, 27], have rapidly emerged over the past few years as a popular approach in modern generative modelling due to their remarkable capabilities in generating complex data, surpassing previous state-of-the-art models, such as Generative Adversarial Networks (GANs) [22] and Variational AutoEncoders (VAEs) [34]. These models are now widely adopted in computer vision and audio generation tasks [35, 11, 47, 2, 32, 3, 53, 51, 72], text generation [41, 70, 45], sequential data modeling [1, 61, 62], reinforcement learning and control [49, 13, 24, 52, 73, 18], as well as life-science [14, 30, 64, 56, 65]. We refer the reader to the survey papers [69, 10] for a more comprehensive exposition of their applications.
The primary goal of SGMs is to generate synthetic data that closely match a target data distribution $\pi_{\mathsf{D}}$, given a sample set. In particular, these models generate approximate data samples from high-dimensional data distributions by combining two diffusion processes, a forward and a backward process in time. The forward process is used to iteratively and smoothly transform samples from the unknown data distribution into (Gaussian) noise, while the associated backward process reverses the noising procedure to generate new samples from the starting unknown data distribution. A key role in these models is played by the score function, i.e. the gradient of the log-density of the solution of the forward process, which appears in the drift of the stochastic differential equation (SDE) associated with the backward process. Since this quantity depends on the unknown data distribution, an estimator of the score must be constructed during the noising step using score-matching techniques [28, 63].
The widespread applicability and success of SGMs have been accompanied by a growing interest in the theoretical understandings of these models, particularly in the convergence analysis under different metrics such as Total Variation (TV) distance, Kullback Leibler (KL) divergence, Wasserstein distance, e.g., [5, 17, 6, 37, 67, 36, 44, 48, 38, 9, 12, 40, 50, 15, 4, 58, 60, 46, 21, 71]. In this work, we provide a non-asymptotic convergence analysis in Wasserstein distance of order two, as this metric is often considered more practical and informative for estimation tasks (see e.g., Equation (5)), and is closely connected to the popular Fréchet Inception Distance (FID) used to assess the quality of images in generative modeling (see, e.g., Section 4). A significant limitation of prior analysis in Wasserstein-2, e.g., [20, 7, 58, 60, 71], is their reliance on strong regularity conditions–such as smoothness or strict log-concavity – of the data distribution and its potential. These assumptions facilitate mathematical tractability but limit the applicability of theoretical results to more general settings, especially when the data distribution is only semiconvex and the potential’s gradient may be discontinuous. The only exception outside the strict log-concavity regime is the recent contribution in [21], where the authors assumes that the data distribution is weakly convex. However, their analysis still requires the potential to be twice continuously differentiable (see, e.g., [21, Proofs of Propositions B.1 and B.2]), and the stepsize of their generative algorithm must be bounded by a quantity inversely proportional to the one-sided Lipschitz constant of the potential (see [21, Equation (30)]). Still, such conditions on $\pi_{\mathsf{D}}$ in existing Wasserstein-2 convergence analysis do not fully reflect the complexity of real-world data, which often exhibit non-smooth or non-log-concave distributions. Therefore, the aim of this work is to address the following fundamental question:
Can Score-based Generative Models be guaranteed to converge in Wasserstein-2 distance when the data distribution is only semiconvex and the potential admits discontinuous gradients?
We provide a positive answer to this question by combining recent findings in non-smooth, non-log-concave sampling, with standard stochastic analysis tools, thereby presenting the first contributions in the Score-based generative modeling literature for non-smooth potentials. We establish explicit, non-asymptotic Wasserstein-2 convergence bounds for SGMs under semiconvexity assumptions on the data distribution, accommodating potentials with discontinuous gradients. This framework covers a variety of practically relevant distributions arising in Bayesian statistical methods, including symmetric modified half-normal distributions, Gaussian mixtures, double-well potentials, and elastic net potentials, all of which satisfy our relaxed assumptions.
In addition, our estimates are explicit and exhibit the best known optimal dependencies in terms of data dimension, i.e., $O(\sqrt{d})$ in Theorem 3.12, and rate of convergence, i.e., $O(\gamma)$ in Theorem 3.14. In contrast to prior works under the same metric [21, 20, 58, 60], our estimates in Theorems 3.12 and Theorem 3.14 are derived without imposing any restrictions on the stepsize of the generative algorithm, making them more suitable to practical implementation. By circumventing the need for strict regularity conditions on the score function and allowing discontinuities in the gradients of the potentials, our work significantly expands the theoretical foundation of SGMs. This advancement not only bridges the gap between empirical success and theoretical guarantees but also opens new avenues for the application of diffusion models to data distributions with non-smooth potentials.
One source of error in the construction of the generative algorithm arises from replacing the initial condition of the backward process with the invariant measure of the forward process. To ensure this error remains small, the drift terms of both SDEs must satisfy, for instance, a monotonicity property with a time-dependent bound that meets an appropriate integrability condition (see, e.g., (19) and (23) below). To address this, we identify a time horizon for the generative algorithm that ensures the paths of the two backward processes become contractive. Notably, the integrability condition on the monotonicity bound depends only on the known constants in Assumption 2, making it significantly easier to verify in practice compared to the analogous condition in [21, Appendix C], which relies on weak convexity constants that are often difficult to estimate.
In conclusion, we present the first explicit, dimension- and parameter-dependent $W_{2}$-convergence guarantees for Score-based Generative models operating on data distributions having potentials with discontinuous gradients. Our results mark an important step forward in the rigorous analysis of SGMs, providing both theoretical insights and practical tools for advancing generative modeling in challenging, non-smooth regimes.
1.1 Notation
[…]
$W_{2}(\mu,\nu)=\left(\inf_{\zeta\in\mathscr{C}(\mu,\nu)}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|x-y|^{2}\ \text{d}\zeta(x,y)\right)^{\frac{1}{2}}.$
For any $x\in A\subseteq\mathbb{R}^{d}$ and any function $U:A\rightarrow\mathbb{R}$, the subdifferential $\partial U(x)$ of $U$ at $x$ is defined as
$\partial U(x)=\left\{p\in\mathbb{R}^{d}:\ \liminf_{y\rightarrow x}\frac{U(y)-U(x)-\langle p,y-x\rangle}{|y-x|}\geq 0\right\}.$ (1)
At the points where $U$ is differentiable, it holds that $\partial U(x)=\{\nabla U(x)\}$. For the sake of presentation, an element of $\partial U(x)$ is denoted by $h(x)$ for any $x\in\mathbb{R}^{d}$.
2 Technical Background for OU-based SGMs
In this section, we provide a brief summary behind the construction of score-based generative models based on diffusion introduced in [57]. The fundamental concept behind SGMs centers around the use of an ergodic (forward) diffusion process to diffuse the unknown data distribution $\pi_{\mathsf{D}}\in\mathscr{P}(\mathbb{R}^{d})$ to a known prior distribution and then learn a backward process to transform the prior to the target distribution $\pi_{\mathsf{D}}$ by estimating the score function of the forward process. In our analysis, we focus on the forward process $(X_{t})_{t\in[0,T]}$ given by an Ornstein-Uhlenbeck (OU) process as chosen in the original paper [57]
$\text{d}X_{t}=-X_{t}\ \text{d}t+\sqrt{2}\ \text{d}B_{t},\quad X_{0}\sim\pi_{\mathsf{D}},$ (2)
where $(B_{t})_{t\in[0,T]}$ is an $d$-dimensional Brownian motion and we assume that $\mathbb{E}[|X_{0}|^{2}]<\infty$. Under mild assumptions on the target data distribution $\pi_{\mathsf{D}}$ [25, 8], the backward process $(Y_{t})_{t\in[0,T]}=(X_{T-t})_{t\in[0,T]}$ is given by
$\text{d}Y_{t}=(Y_{t}+2\nabla\log p_{T-t}(Y_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}\sim\mathscr{L}(X_{T}),$ (3)
where $\{p_{t}\}_{t\in[0,T]}$ is the family of densities of $\{\mathscr{L}(X_{t})\}_{t\in(0,T]}$ with respect to the Lebesgue measure, $\bar{B}_{t}$ is an another Brownian motion independent of $B_{t}$ in (2) defined on $(\Omega,\mathscr{F},\mathbb{P})$. However, the initial distribution in practice is the invariant measure of the forward process, which, in this case, is a standard $d$-dimensional Gaussian distribution. Therefore, the backward process (3) becomes
$\text{d}\widetilde{Y}_{t}=(\widetilde{Y}_{t}+2\ \nabla\log p_{T-t}(\widetilde{Y}_{t}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad\widetilde{Y}_{0}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}).$ (4)
Since $\pi_{\mathsf{D}}$ is unknown, the score function $\nabla\log p_{t}$ in (3) cannot be computed exactly. To address this issue, an estimator $s(\cdot,\theta^{*},\cdot)$ is learned based on a family of functions $s:[0,T]\times\mathbb{R}^{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ parametrized in $\theta$, aiming at approximating the score of the ergodic (forward) diffusion process over a fixed time window $[0,T]$. In practice, the functions $s$ are neural networks and in particular cases, e.g. the motivating example in [7, Section 3.1], the functions $s$ can be wisely designed. The optimal value $\theta^{*}$ of the parameter $\theta$ is determined by optimizing the following score-matching objective
$\begin{split}\mathbb{R}^{d}\ni\theta\mapsto\mathbb{E}\left[\int_{0}^{T}|\nabla\log p_{t}(X_{t})-s(t,\theta,X_{t})|^{2}\ \mathrm{d}t\right].\end{split}$ (5)
An explicit expression of the stochastic gradient of (5) derived via denoising score matching [63] is provided in [7, Equation (8), Section 2]. Following [7, Section 2], we introduce an auxiliary process $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ containing the approximating function $s$ depending on the (random) estimator of $\theta^{*}$ denoted by $\hat{\theta}$, for $t\in[0,T]$,
$\text{d}Y_{t}^{\text{aux}}=(Y_{t}^{\text{aux}}+2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\quad Y_{0}^{\text{aux}}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}).$ (6)
The process (6) connects the backward process (4) and the numerical scheme (8), which facilitates the analysis of the convergence of the diffusion model (see the upper bounds in Appendix C involving $Y_{t}^{\text{aux}}$ for more details). Let the step size $\gamma_{k}=\gamma\in(0,1)$ for each $k=0,\dots,K$, where $K\in\mathbb{N}$. The discrete process $(Y_{k}^{\text{EM}})_{k\in\{0,\dots,K+1\}}$ of the Euler–Maruyama approximation of (6) is given, for any $k\in\{0,\dots,K\}$, as follows
$Y_{k+1}^{\text{EM}}=Y_{k}^{\text{EM}}+\gamma(Y_{k}^{\text{EM}}+2\ s(T-t_{k},\hat{\theta},Y_{k}^{\text{EM}}))+\sqrt{2\gamma}\ \bar{Z}_{k+1},\quad Y_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}),$ (7)
where $\{\bar{Z}_{k}\}_{k\in\{0,\dots,K+1\}}$ is a sequence of independent $d$-dimensional Gaussian random variables with zero mean and identity covariance matrix. The continuous-time interpolation of (7), for $t\in[0,T]$, is given by
$\mathrm{d}\widehat{Y}_{t}^{\text{EM}}=(\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}+2\ s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}))\ \text{d}t+\sqrt{2}\ \text{d}\bar{B}_{t},\qquad\widehat{Y}_{0}^{\text{EM}}\sim\pi_{\infty}=\mathscr{N}(0,I_{d}),$ (8)
where $\mathscr{L}(\widehat{Y}_{k}^{\text{EM}})=\mathscr{L}(Y^{\text{EM}}_{k})$ at grid points for each $k\in\{0,\dots,K+1\}$.
3 Wasserstein Convergence Analysis for SGMs
In this section, we provide the full non-asymptotic estimates in Wasserstein distance of order two between the target data distribution $\pi_{\mathsf{D}}$ and the generative distribution of the diffusion model under the assumptions stated below. As explained in [7, Section 2 and Appendix A]), it could be necessary to restrict $t\in[\epsilon,T]$ for $\epsilon\in(0,1)$ in (5) to account for numerical instability issues for training and sampling at $t=0$ as observed in practice in [57, Appendix C] and for the possibility that the integral of the score function in (5) may diverge when $t=0$. Therefore, we truncate the integration in the backward diffusion at $T-\epsilon$ and run the process $(Y_{t})_{t\in[0,T-\epsilon]}$.
3.1 Assumptions
We begin by stating the main assumptions of our setting. The optimization problem in (5) can be solved using algorithms such as ADAM [33], stochastic gradient descent [29], Stochastic Gradient Langevin Dynamics [7, Section 3.1], and TheoPouLa [43], provided they satisfy the following assumption.
Assumption 1.
Let $\theta^{*}$ be a minimiser [^1] The score-matching optimization problem (5) is not necessarily (strongly) convex. of (5) and let $\hat{\theta}$ be the (random) estimator of $\theta^{*}$ obtained through some approximation procedure such that $\mathbb{E}[|\hat{\theta}|^{2}]<\infty$. There exists $\widetilde{\varepsilon}_{\text{AL}}>0$ such that
$\mathbb{E}[|\hat{\theta}-\theta^{*}|^{2}]<\widetilde{\varepsilon}_{\text{AL}}.$
Remark 3.1.
As a consequence of Assumption 1, one obtains $\mathbb{E}[|\hat{\theta}|^{2}]<2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2}$.
Recall that for any $x\in\mathbb{R}^{d}$, $h(x)$ denotes an element of the subdifferential $\partial U(x)$ defined in (1).
Assumption 2.
The data distribution $\pi_{\mathsf{D}}$ has a finite second moment and it is absolutely continuous with respect to the Lebesgue measure with $\pi_{\mathsf{D}}(\mathrm{d}x)=\exp(-U(x))\ \mathrm{d}x$ for some $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover,
  1. (i)
    The potential $U$ is continuous and its gradient exists almost everywhere.
  2. (ii)
    The potential $U$ is $K$-semiconvex. That is, there exists $K,R\geq 0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq-K|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|<R,$
    or equivalently $U+\frac{K}{2}|\cdot|^{2}$ is convex.
  3. (iii)
    The potential $U$ is $\mu$-strongly convex at infinity. That is, there exists $\mu>0$ and $R\geq 0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq\mu|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|\geq R.$ (9)
Remark 3.2.
As a consequence of Proposition A.1, due to in [15, Proposition 3.1], and Assumption 2-(i), $\nabla\log p_{t}(x)$ is continuous for $t\in(0,T]$ and $x\in\mathbb{R}^{d}$. Moreover, Assumption 2 implies that the processes in (3) and (4) have a unique strong solution.
Next, we consider the following assumption on the approximating function $s$.
Assumption 3.a.
The function $s:[0,T]\times\mathbb{R}^{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is continuously differentiable in $x\in\mathbb{R}^{d}$. Let $D_{1}:\mathbb{R}^{M}\times\mathbb{R}^{M}\rightarrow\mathbb{R}_{+}$, $D_{2}:[0,T]\times[0,T]\rightarrow\mathbb{R}_{+}$ and $D_{3}:[0,T]\times[0,T]\rightarrow\mathbb{R}_{+}$ be such that $\int_{\epsilon}^{T}\int_{\epsilon}^{T}D_{2}(t,\bar{t})\ \text{d}t\ \text{d}\bar{t}<\infty$ and $\int_{\epsilon}^{T}\int_{\epsilon}^{T}D_{3}(t,\bar{t})\ \text{d}t\ \text{d}\bar{t}<\infty$. For $\alpha\in\left[\frac{1}{2},1\right]$ and for all $t,\bar{t}\in[0,T]$, $x,\bar{x}\in\mathbb{R}^{d}$, and $\theta,\bar{\theta}\in\mathbb{R}^{M}$, we have that
$\begin{split}|s(t,\theta,x)-s(\bar{t},\bar{\theta},\bar{x})|&\leq D_{1}(\theta,\bar{\theta})|t-\bar{t}|^{\alpha}+D_{2}(t,\bar{t})|\theta-\bar{\theta}|+D_{3}(t,\bar{t})|x-\bar{x}|,\end{split}$
where $D_{1}$, $D_{2}$ and $D_{3}$ have the following growth in each variable: i.e. there exist $\mathsf{K}_{1}$, $\mathsf{K}_{2}$, and $\mathsf{K}_{3}>0$ such that for each $t,\bar{t}\in[0,T]$ and $\theta,\bar{\theta}\in\mathbb{R}^{M}$,
$\begin{split}|D_{1}(\theta,\bar{\theta})|&\leq\mathsf{K}_{1}(1+|\theta|+|\bar{\theta}|),\qquad|D_{2}(t,\bar{t})|\leq\mathsf{K}_{2}(1+|t|^{\alpha}+|\bar{t}|^{\alpha}),\\ |D_{3}(t,\bar{t})|&\leq\mathsf{K}_{3}(1+|t|^{\alpha}+|\bar{t}|^{\alpha}).\end{split}$
Remark 3.3.
Assumption 3.a implies that the process in (6), (7), and (8) have a unique strong solution. We refer to [7, Remark 6] for discussion regarding the practical justification of this assumption when it is applied to neural networks.
By adding a further condition on the gradient of $s$ in Assumption 3.a we are able to achieve the optimal rate of convergence in Theorem 3.14 below.
Assumption 3.b.
Let $s$ be as in Assumption 3.a and there exists $\mathsf{K}_{4}>0$ such that, for all $x,\bar{x}\in\mathbb{R}^{d}$ and for any $k=1,\dots d$,
$|\nabla_{x}s^{(k)}(t,\theta,x)-\nabla_{\bar{x}}s^{(k)}(t,\theta,\bar{x})|\leq\mathsf{K}_{4}(1+2|t|^{\alpha})|x-\bar{x}|.$
For the following assumption on the score approximation, we let $\hat{\theta}$ be as in Assumption (1) and we let $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ be the auxiliary process defined in (6).
Assumption 4.
There exists $\varepsilon_{\text{SN}}>0$ such that
$\mathbb{E}\int_{0}^{T-\epsilon}|\nabla\log p_{T-r}(Y_{r}^{\text{aux}})-s(T-r,\hat{\theta},Y_{r}^{\text{aux}})|^{2}\ \text{d}r<\varepsilon_{\text{SN}}.$ (10)
Remark 3.4.
We refer to [7, Remark 7, 8, and 9] for discussion regarding the theoretical and practical soundedness of Assumption 4, which is a now a standard assumption considered in the literature, see e.g. [7, 20, 21, 58].
3.2 Assumption 2 and Weak Convexity of the Data Distribution
We show that Assumption 2-(ii) and Assumption 2-(iii) are related to the notion of weak convexity in the sense made precise in Proposition 3.7 below. We start by introducing the definition of weak convexity for subgradients.
Definition 3.5.
The potential $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is weakly convex if its weak convexity profile $\kappa_{U}:[0,\infty)\rightarrow\mathbb{R}$ defined as
$\kappa_{U}(r)=\inf_{x,\bar{x}\in\mathbb{R}^{M}:\ |x-\bar{x}|=r}\left\{\frac{\langle\partial U(x)-\partial U(\bar{x}),x-\bar{x}\rangle}{|x-\bar{x}|^{2}}\right\}$ (11)
satisfies
$\kappa_{U}(r)\geq\beta-r^{-1}f_{L}(r),\quad\text{for all}\ \ r>0.$ (12)
for some constants $\beta,L>0$, where the function $f_{L}:[0,\infty]\rightarrow[0,\infty]$ is defined as
$f_{L}(r)=2L^{1/2}\tanh((rL^{1/2})/2).$ (13)
[…]
Lemma 3.6.
[16, Modification of Lemma 5.9] Assume that U is weakly convex as in Definition 3.5 and fix $t\in[0,T]$. Then, the function $x\mapsto-\log p_{t}(x)$ is weakly convex with weak convexity profile $\kappa_{-\log p_{t}(x)}$ satisfying
$\kappa_{-\log p_{t}}(r)\geq\frac{\beta}{\beta+(1-\beta)e^{-2t}}-\frac{e^{-t}}{\beta+(1-\beta)e^{-2t}}\frac{1}{r}f_{L}\left(\frac{e^{-t}}{\beta+(1-\beta)e^{-2t}}r\right).$
In particular, the score function satisfies
$\langle\nabla\log p_{t}(x)-\nabla\log p_{t}(\bar{x}),x-\bar{x}\rangle\leq-\widehat{C}_{t}|x-\bar{x}|^{2},\quad\text{for}\ x,\bar{x}\in\mathbb{R}^{\text{d}}$ (14)
with
$\widehat{C}_{t}=\frac{\beta}{\beta+(1-\beta)e^{-2t}}-\frac{e^{-2t}}{(\beta+(1-\beta)e^{-2t})^{2}}L.$ (15)
An overview of the proof of Proposition 3.7 below can be found in Appendix B.
Proposition 3.7.
Let the data distribution $\pi_{\mathsf{D}}$ be in Assumption 2 and let $f_{L}$ be as in (13). Then
$\begin{split}\kappa_{U}(r)\geq\mu-r^{-1}f_{L}(r),\quad\text{for all}\ r>0,\end{split}$ (16)
where $\mu>0$ is the strong convexity at infinity constant from Assumption 2-(iii). Conversely, if $U$ is weakly convex as in Definition 3.5 with lower bound (16) for some known constants $\mu$ and $L>0$, then
  1. 1.
    The potential $U$ is $\widetilde{\mu}$-strongly convex at infinity with $\widetilde{\mu}:=\mu-R^{-1}f_{L}(R)>0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$, we have
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq\widetilde{\mu}|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|\geq R,$ (17)
    which holds for all $R>0$ when $\mu>L$ and for $R\geq R_{0}=\frac{2z_{0}}{L^{1/2}}$ with $z_{0}$ being the solution of (43) when $\mu\leq L$.
  2. 2.
    The potential $U$ is $K$-semiconvex, such that there exists $K,R\geq 0$ for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq-K|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|\leq R.$ (18)
As a consequence of Proposition 3.7 and Lemma 3.6, one obtains the explicit form of $\widehat{C}_{t}$ in (14) in our setting, which is given in the following corollary.
Corollary 3.8.
Let $U$ be $K$-semiconvex as in Assumption 2-(ii) and be $\mu$-strongly convex at infinity as in Assumption 2-(iii) and fix $t\in[0,T]$. Then
$\langle\nabla\log p_{t}(x)-\nabla\log p_{t}(\bar{x}),x-\bar{x}\rangle\leq-\beta_{t}^{\text{OS}}|x-\bar{x}|^{2},\qquad\text{for}\quad x,\bar{x}\in\mathbb{R}^{d},$ (19)
where
$\beta^{\text{OS}}_{t}=\frac{\mu}{\mu+(1-\mu)e^{-2t}}-\frac{e^{-2t}}{(\mu+(1-\mu)e^{-2t})^{2}}L,$ (20)
for some $L>0$ satisfying (16).
Remark 3.9.
By Corollary 3.8 and the proof of Proposition 3.7, we have
$\lim_{t\rightarrow 0}\beta^{\text{OS}}_{t}=\mu-L<-K.$ (21)
[…]
For this reason, we use the constant $K+\mu$ as a proxy of the constant $L$ and replace (20) with the following monotonicity bound
$\beta^{\text{OS},K,\mu}_{t}=\frac{\mu}{\mu+(1-\mu)e^{-2t}}-\frac{e^{-2t}}{(\mu+(1-\mu)e^{-2t})^{2}}(K+\mu).$ (22)
Moreover, it holds that
$\lim_{t\rightarrow 0}\beta^{\text{OS},K,\mu}_{t}=-K,$
and
$\lim_{t\rightarrow\infty}\beta^{\text{OS}}_{t}=\lim_{t\rightarrow\infty}\beta^{\text{OS},K,\mu}_{t}=1,$
which is consistent with $\pi_{\infty}\sim\mathscr{N}(0,I_{d})$, the invariant distribution of the OU process.
Using the explicit expression of (22), we are able to find a time for which the integral of the monotonicity bound [^3] Note that $\beta^{\text{OS},K,\mu}_{t}$ is a function of time. $\beta^{\text{OS},K,\mu}_{t}$ is positive. The proof of the following result is postponed to Appendix B.
Proposition 3.10.
Let $\mu>0$ and $K\geq 0$. The time integral of $\beta^{\text{OS},K,\mu}_{t}$ from Remark 3.9 is
$\begin{split}B(t,0,\mu,K)&=\int_{0}^{t}\left(\frac{\mu}{\mu+(1-\mu)e^{-2s}}-\frac{e^{-2s}}{(\mu+(1-\mu)e^{-2s})^{2}}(K+\mu)\right)\ \mathrm{d}s\\ &=\frac{1}{2}\left[\log\left(\mu(e^{2t}-1)+1\right)+\left(\frac{K}{\mu}+1\right)\left(\frac{1}{\mu(e^{2t}-1)+1}-1\right)\right]>0,\end{split}$ (23)
when $t>t^{\star}>\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)$ with $t^{\star}:=\inf\left\{t>0:B(t,0,\mu,K)>0\right\}$.
Remark 3.11.
If we consider the case when $K=0$ in Assumption 2-(ii), then (23) is satisfied for all $t>0$.
3.3 Main Results - Optimal Data Dimensional Dependence and Rate of Convergence
The main results are stated as follows. An overview of their proofs can be found in Appendix C.
Theorem 3.12.
Let Assumptions 1, 2, 3.a and 4 hold. Then, there exist constants $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}>0$ such that for any $T>0$ and $\gamma,\epsilon\in(0,1)$,
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+C_{4}(T,\epsilon)\gamma^{1/2},\end{split}$ (24)
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 2 (Appendix E), $\beta^{\text{OS},K,\mu}_{t}$ is defined in (22), and its integral is computed in Proposition 3.10. In addition, the result in (24) implies that for any $\delta>0$, if we choose $0<\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 2, then
$W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\leavevmode\nobreak\ \delta.$
Remark 3.13.
The constant $C_{4}(T,\epsilon)$ in the error bound in (24) contains the optimal dependence of the data dimension, i.e. $O(\sqrt{d})$, which has been found under the more strict assumption of strong-log concavity of $\pi_{\mathsf{D}}$ in [7, Theorem 1 and Remark 12]. However, the optimal dependence of the dimension is achieved at the expenses of a worst rate of covergence of order $1/2$.
The optimal rate of convergence of order $\alpha\in[\frac{1}{2},1]$ for the Euler or Milstein scheme of SDEs with constant diffusion coefficients can be achieved in Theorem 3.12 if we require that $\mathbb{E}[|\hat{\theta}|^{4}]<\infty$ in Assumption 1 and replace Assumption 3.a with Assumption 3.b as stated in Theorem 3.14 below.
Theorem 3.14.
Let Assumptions 1, 2, 3.b and 4 hold, and assume that $\mathbb{E}[|\hat{\theta}|^{4}]<\infty$ Then, there exist constants $C_{1}$, $C_{2}$, $C_{3}$ and $\widetilde{C}_{4}>0$ such that for any $T>0$ and $\gamma,\epsilon\in(0,1)$,
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+\widetilde{C}_{4}(T,\epsilon)\gamma^{\alpha},\end{split}$ (25)
where $C_{1}$, $C_{2}$, $C_{3}$ and $\widetilde{C}_{4}$ are given explicitly in Table 2 (Appendix E), $\beta^{\text{OS},K,\mu}_{t}$ is defined in (22), and its integral is computed in Proposition 3.10. In addition, the result in (25) implies that for any $\delta>0$, if we choose $0<\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\widetilde{\gamma}_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\widetilde{\gamma}_{\delta}$ given in Table 2, then
$W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\leavevmode\nobreak\ \delta.$ (26)
Remark 3.15.
In the explicit expression of $\widetilde{C}_{4}$ in Table 2 (Appendix E), the dependence of the data dimension is $O(d)$ due to numerical techniques from [7, Proof of Theorem 10] used in the proof of Theorem 3.14 to achieve the optimal rate of convergence of order $\alpha\in[\frac{1}{2},1]$.
3.4 Examples of potentials satisfying by Assumption 2
We present several examples to demonstrate the wide applicability of our Assumption 2 to a broad class of data distributions, some of which are not covered by previous results in Wasserstein distance of order two [21, 7, 20, 58, 60, 71].
3.4.1 Symmetric modified half-normal distribution
We consider the case of a one-dimensional symmetric modified half-normal distribution
$\begin{split}\pi_{\mathsf{D}}(\mathrm{d}x)=\frac{\sqrt{\xi}\exp\left(-\xi x^{2}-|x|\right)}{\Psi\left(\frac{1}{2},\frac{-1}{\sqrt{\xi}}\right)}\ \mathrm{d}x,\quad x\in\mathbb{R},\end{split}$ (27)
for some unknown $\xi>0$ and normalizing constant
$\begin{split}\Psi\left(\frac{1}{2},\frac{-1}{\sqrt{\xi}}\right)&:=\sum_{n=0}^{\infty}\frac{\Gamma\left(\frac{1}{2}+\frac{n}{2}\right)}{\Gamma(n)}\frac{(-1)^{n}\xi^{-n/2}}{n!},\end{split}$
where $\Gamma(n)$ is the Gamma function. We refer the reader to Appendix D for additional details about the derivation of (27). As highlighted in [59, Section 2], the modified half-normal distribution appears in several Bayesian statistical methods as a posterior distribution to sample from in Bayesian Binary regression, analysis of directional data, and Bayesian graphical models.
Assumption 2-(i) is satisfied for $U(x)=\xi x^{2}+|x|$. In addition, we have, for all $x,\bar{x}\in\mathbb{R}$
$\begin{split}\langle h(x)-h(\bar{x}),x-\bar{x}\rangle&=2\left(\xi|x-\bar{x}|^{2}+(x-\bar{x})\mathbbm{1}_{x>0,\ \bar{x}<0}-(x-\bar{x})\mathbbm{1}_{x<0,\ \bar{x}>0}\right)\\ &\geq 2\xi|x-\bar{x}|^{2},\end{split}$ (28)
which shows that Assumption 2-(ii) is verified for any $K\geq 0$, and Assumption 2-(iii) is verified for $\mu=2\xi$. Therefore, we can conclude that (27) satisfies Assumption 2.
3.4.2 Multidimensional Gaussian mixture distribution
We consider a multidimensional Gaussian mixture data distribution with unknown mean and variance, i.e.
$\pi_{\mathsf{D}}(\mathrm{d}x)=\sum_{j=1}^{J}\widetilde{\xi}_{j}\frac{1}{(2\pi\sigma_{j}^{2})^{d/2}}\exp\left(-\frac{|x-\eta_{j}|^{2}}{2\sigma_{j}^{2}}\right)\ \mathrm{d}x,\quad x\in\mathbb{R}^{d},$ (29)
with $\sigma_{j}>0$, $\eta_{j}\in\mathbb{R}^{d}$, and $\widetilde{\xi}_{j}\in[0,1]$ for $j\in\{1,\dots,J\}$ such that $\sum_{j=1}^{J}\widetilde{\xi}_{j}=1$. The authors in [21, Appendix A] show that the score function of (29) is Lipschitz continuous and $-\log\pi_{\mathsf{D}}$ is weakly convex. Therefore, Assumption 2 is satisfied. In addition, the distribution (29) covers also case of the double-well potential:
$U(x)=x^{4}-|x|^{2},\quad x\in\mathbb{R}^{d},$ (30)
which is $2$-semiconvex and strongly convex at infinity.
3.4.3 Multi-dimensional Potentials
Similarly as in Section 3.4.1, one can proves that the elastic net potential:
$U(x)=|x|^{2}+\sum_{i=1}^{d}|x_{i}|,\quad x\in\mathbb{R}^{d},$ (31)
satisfies Assumption 2. Moreover, the following potential
$U(x)=\max\left\{|x|,|x|^{2}\right\},\qquad x\in\mathbb{R}^{d},$ (32)
verifies Assumption 2 with $K=0$, $R=1$, and $\mu=2$ as well as the following non-convex potential presented in [31, Example 4.2]:
$U(x)=\max\left\{|x|,|x|^{2}\right\}-\frac{1}{2}|x|^{2},\qquad x\in\mathbb{R}^{d}.$ (33)
4 Related Work and Comparison
In recent years, there has been a rapidly expanding body of research on the convergence theory of Score-based Generative Models. Existing works for convergence bounds can be divided into two main approaches, depending on the divergence or distance used.
The first approach focuses on $\alpha$-divergences, particularly the Kullback–Leibler (KL) divergence and Total Variation (TV) distance (e.g., [4, 15, 68, 39, 5, 17, 37, 40, 38, 9, 12, 48, 42, 67]), which are the vast majority of the results available in the literature. Crucially, bounds on KL divergence imply bounds on TV distance via Pinsker’s inequality, strengthening their wide applicability. We provide a brief and selective overview of some of the findings following this first approach. The results in TV distance in [37] and in KL divergence [68] assumed that the data distribution satisfies a logarithmic Sobolev inequality and the score function was Lipschitz resulting in convergence bounds characterized by polynomial complexity. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in [12] managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time instead of the whole trajectory, the authors in [9, Theorem 2.5] managed to show, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. In the work by [4], the authors provide convergence bounds in KL divergence that are linear in the data dimension, up to logarithmic factors, using early stopping and assuming finite second moments of the data distribution. However, both results in [9, Theorem 2.5] and [4, Theorem 1 and Corollary 1] still require the uniqueness of solutions for the backward SDE (3), and therefore additional assumptions on the score function are needed. We refer the reader to [7, Section 4.2] for additional details about this point. Assuming the finiteness of the second moment of the data distribution and using an EI discretization scheme with constant and exponentially decreasing step sizes, the authors in [15, Corollary 2.4] derive a KL divergence bound with early stopping, which scales linearly in the data dimension up to logarithmic factors. Bounds in KL without early stopping have been derived in [15] for data distributions with finite Fisher information with respect to the standard Gaussian distribution. We note that this assumption stated in [15, Assumption H2] still requires the potential $U\in C^{1}(\mathbb{R}^{d})$. The bounds provided in [15, Theorem 2.1 and 2.2] scale linearly in the Fisher information when an EI discretization scheme with constant step size is used and logarithmically in the Fisher information when an exponential-then-constant step size [15, Theorem 2.3] is chosen.
The second approach focuses on convergence bounds in Wasserstein distance, a metric which is often considered more practical and informative for estimation tasks. We can relate results following this approach with the results of the first approach only when $\pi_{\mathsf{D}}$ is a strongly log-concave distribution. In this case, $W_{2}$-bounds in terms of KL divergence follow from an extension of Talagrand’s inequality [23, Corollary 7.2]. However, for two general data distributions, there is no known relationship between their KL divergence and their $W_{2}$. Therefore, we cannot compare our findings in Therem 3.12 and Theorem 3.14 with the results derived following the first approach. One line of work within the second approach assumes (at least) strong log-concavity of the data distribution [20, 7, 60, 58, 71]. Under this (strict) assumption, [7, Remark 12] achieved optimal data dimensional dependence, i.e. reaching $O(\sqrt{d})$. The recent bound in [21, Theorem D.1] scales linearly in the data dimension while relaxing the strong log-concavity assumption on $\pi_{\mathsf{D}}$ to weakly log-concavity, but still requiring that the potential $\nabla^{2}U$ exists (see, e.g., [21, Proof of Proposition B.1 and B.2]). Our Assumption 2 is much weaker than this requirement and it allows to consider the case of potentials with discontinuous gradients covering a wider range of distributions as outlined in Section 3.4. Another line of work following this approach focuses on specific structural assumptions of the data distribution. For instance, convergence bounds in Wasserstein distance of order one with exponential dependence on the problem parameters have been obtained in [6] under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Under the same metric, the authors in [46] provide a convergence analysis when the data distribution is defined on a torus. We summarize in Table 1 and the best results obtained in $W_{2}$, i.e. [7, 21] and compare with our best result, which scale polynomially in the data dimension, i.e. $O(\sqrt{d})$ (Theorem 3.12).
Table 1: Summary of previous bounds for $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ and our result in Theorem 3.12. All the bounds assume that $\pi_{\text{D}}(\mathrm{d}x)\propto e^{-U(x)}\mathrm{d}x$ has finite second moments.
Assumption on $\pi_{\text{D}}$ Error bound Reference
$U$ strongly convex, $\nabla\log p_{t}(0)\in L^{2}([\epsilon,T])$, and Assumption 4 $O(\sqrt{d})\sqrt{\epsilon}+O(\sqrt{d})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+O(e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)})\sqrt{\varepsilon_{\text{SN}}}+O(\sqrt{d}e^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})\gamma^{1/2}$, with $\widehat{L}_{\text{MO}}>0$ lower bound of the strongly convex constant of $U$, see e.g., [7, Remark 4]. [7, Remark 12]
$U\in C^{2}(\mathbb{R}^{d})$, weakly convex, and Assumption 4 $e^{(2L_{U}+5)\eta(\beta,L,\gamma)}[e^{-T}W_{2}(\pi_{\mathsf{D}},\pi_{\infty})+4\varepsilon_{\text{SN}}(T-\eta(\beta,L,0))+\sqrt{2\gamma}(4L_{U}d+6d+\sqrt{d+\mathbb{E}[|X_{0}|^{2}]})(T-\eta(\beta,L,0))]$, with $L_{U}\geq 0$ one-sided Lipschitz constant for $\nabla U$, see e.g., [21, Assumption H1], $\eta(\beta,L,\gamma)$ defined in [21, Equation (29)], and $\gamma<2/(2L_{U}+5)^{2}$ [21, Theorem D.1]
Assumption 2 and Assumption 4 $O(\sqrt{d})\sqrt{\epsilon}+O(\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}+O(e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t})\sqrt{\varepsilon_{\text{SN}}}+O(\sqrt{d}e^{T^{2\alpha+1}}T^{3\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})\gamma^{1/2}$ Theorem 3.12
We close this section by briefly commenting on the choice of deriving our results in Wasserstein distance of order two. Beyond its theoretical relevance, this choice is motivated by practical considerations in generative modeling. First, the Wasserstein distance is often regarded as a more informative and robust metric for estimation tasks. Second, a widely used performance metric for evaluating the quality of images produced by generative models is the Fréchet Inception Distance (FID) [26], which measures the Fréchet distance between the distributions of generated and real samples, assuming Gaussian distributions. In particular, this Fréchet distance is equivalent to the Wasserstein-2 distance. Thus, providing convergence results under the Wasserstein-2 metric enhances the practical relevance of our theoretical findings.
Acknowledgments.
This work was supported by Innovate UK [grant number 10081810]. This work has received funding from the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828).
Appendix A Regularity of the score function
We recall the following result to justify the smoothness of the map
$(0,T]\times\mathbb{R}^{d}\ni(t,x)\mapsto p_{t}(x)\in\mathbb{R}_{+},$ (34)
where $p_{t}$ density of the forward process defined in Section 2.
Proposition A.1.
[15, Proposition 3.1] Let $\pi_{\mathsf{D}}$ be absolutely continuous with respect to the Lebesgue measure, and denote its density by $p_{0}$. The map defined in (34) is positive and solution of the following Fokker–Planck equation on $(0,T]\times\mathbb{R}^{d}$:
$\partial_{t}p_{t}(x)-\text{div}(x\ p_{t})-\Delta p_{t}(x)=0,\quad\text{for}\ (t,x)\in(0,T]\times\mathbb{R}^{d}.$
Moreover, it belongs to $C^{1,2}((0,T]\times\mathbb{R}^{d})$; i.e. for any $t\in(0,T)$, $x\mapsto p_{t}(x)$ is twice continuously differentiable, and for any $x\in\mathbb{R}^{d}$, $t\mapsto p_{t}(x)$ is continuously differentiable on $(0,T]$.
Appendix B Further Details on Assumption 2 and Weak Convexity of the Data Distribution
We provide the proofs of Section 3.2.
Proof of Proposition 3.7.
We begin by considering that $\pi_{\mathsf{D}}$ satisfies Assumption 2. Recall that $f_{L}$ is defined as in (13). Note that $r\mapsto r^{-1}f_{L}(r)$ is non-increasing on $(0,\infty)$ and $f_{L}^{{}^{\prime}}(0)=L>r^{-1}f_{L}(r)$ for $r\in(0,R]$. We look for $L>0$ satisfying
$\inf_{r\in(0,R]}r^{-1}f_{L}(r)=R^{-1}f_{L}(R)=2R^{-1}L^{1/2}\tanh((RL^{1/2})/2)=K+\mu.$ (35)
Equivalently, we look for $x=L^{1/2}R/2>0$ such that
$x\tanh(x)=\frac{K+\mu}{4}R^{2},\quad\text{subject to}\quad x>\frac{\sqrt{K+\mu}}{2}R,$ (36)
so as $L>K+\mu$. Note that $\tanh(x)\leq x$ for all $x\geq 0$ Therefore, if we choose $x=\frac{\sqrt{K+\mu}}{2}R$, then
$\frac{\sqrt{K+\mu}}{2}R\tanh\left(\frac{\sqrt{K+\mu}}{2}R\right)\leq\frac{K+\mu}{4}R^{2}.$ (37)
Using (37) and $\lim_{x\uparrow\infty}x\tanh(x)=\infty$, we deduce that there exists $x^{\star}>0$ such that
$x^{\star}\tanh(x^{\star})=\frac{K+\mu}{4}R^{2},$ (38)
with $x^{\star}>\frac{\sqrt{K+\mu}}{2}R$, since $x\mapsto x\tanh(x)$ is non-decreasing on $(0,\infty)$. By Assumption 2 and (35), we have
$\begin{split}k_{U}(r)&\geq\mu-(K+\mu)\\ &\geq\mu-r^{-1}f_{L}(r),\qquad\text{for}\ r\leq R.\end{split}$ (39)
Moreover,
$\begin{split}k_{U}(r)&\geq\mu\\ &\geq\mu-r^{-1}f_{L}(r),\qquad\text{for}\ r>R,\end{split}$
where it is used that $r^{-1}f_{L}(r)>0$ for all $r>0$. This proves the first part of the statement in Proposition 3.7, i.e. the lower bound (16).
Conversely, assume that $U$ is weakly convex as in Definition 3.5 with lower bound (16) for some known constants $\mu$ and $L>0$. We look for $R$ such that
$\begin{split}\kappa_{U}(r)&\geq\mu-r^{-1}f_{L}(r)\\ &\geq\mu-R^{-1}f_{L}(R)\\ &>0,\qquad\qquad\qquad\qquad\forall\ r>R,\end{split}$ (40)
where it is used that $r^{-1}f_{L}(r)$ is decreasing on $(0,\infty)$. Let $\widetilde{\mu}:=\mu-R^{-1}f_{L}(R)$, so (40) becomes $\kappa_{U}(r)\geq\widetilde{\mu}>0$, for all $r>R$. One notes that
$\begin{split}\widetilde{\mu}&=\mu-L\frac{\tanh((RL^{1/2})/2)}{(RL^{1/2})/2}>0.\end{split}$ (41)
If $\mu>L$, (41) is satisfied for all $R>0$. If $\mu\leq L$, (41) holds for $R\geq R_{0}$, where $R_{0}$ is the unique solution to
$\mu=\frac{2L^{1/2}}{R}\tanh\left(\frac{RL^{1/2}}{2}\right).$ (42)
Let $z=\frac{RL^{1/2}}{2}$, then $R_{0}=\frac{2z_{0}}{L^{1/2}}$, where $z_{0}$ solves
$\frac{\tanh(z)}{z}=\frac{\mu}{L}.$ (43)
Since $\frac{\tanh(z)}{z}$ monotonically decreases from $1$ to 00 as $z$ increases, a unique $z_{0}>0$ solving (43) exists for $\mu<L$. Therefore, (40) is satisfied for $R\geq R_{0}=\frac{2z_{0}}{L^{1/2}}$. This proves that $U$ is $\widetilde{\mu}$-strongly convex at infinity, and therefore (17). Using the assumption that $U$ is weakly convex as in Definition 3.5, one obtains that
$\begin{split}\kappa_{U}(r)&\geq\mu-r^{-1}f_{L}(r)\\ &\geq\mu-L,\qquad\qquad\text{for}\quad r\leq R.\end{split}$ (44)
We distinguish two cases for the lower bound in (44). If $\mu>L$, then $\kappa_{U}(r)\geq-K$ for $r\leq R$ for all $R>0$ and $K\geq 0$. If $\mu\leq L$, then, by setting $K=L-\mu$ in (43), we have $\kappa_{U}(r)\geq-K$ for $r\leq R$ for all $R>0$. This proves that $U$ is $K$-semiconvex, and therefore (18). This concludes the proof for the second part of the statement in Proposition 3.7.
Proof of Proposition 3.10.
We look for $t^{\star}$ satisfying
$\begin{split}B(t^{\star},0,\mu,K)=\frac{1}{2}\left[\log\left(\mu(e^{2t^{\star}}-1)+1\right)+\left(\frac{K}{\mu}+1\right)\left(\frac{1}{\mu(e^{2t^{\star}}-1)+1}-1\right)\right]>0.\end{split}$ (45)
Equivalently, we look for $x:=e^{2t^{\star}}-1$ such that
$\begin{split}g(\mu x+1)=\log(\mu x+1)-\left(\frac{K}{\mu}+1\right)\frac{\mu x}{\mu x+1}>0.\end{split}$ (46)
Note that (46) is satisfied for all $x>0$ when $K=0$. In addition, we have
$\begin{split}\lim_{x\rightarrow 0+}g(\mu x+1)&=0,\\ \lim_{x\rightarrow+\infty}g(\mu x+1)&=\infty,\end{split}$ (47)
and
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}x}g(\mu x+1)&=\frac{\mu}{\mu x+1}-\frac{K+\mu}{(\mu x+1)^{2}}\geq 0\quad\text{when}\quad x\geq\frac{K}{\mu^{2}}.\\ \frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}g(\mu x+1)&=-\frac{\mu^{2}}{(\mu x+1)^{2}}+\frac{2(K+\mu)\mu}{(\mu x+1)^{3}}\geq 0\quad\text{when}\quad x\leq\frac{2K}{\mu^{2}}+\frac{1}{\mu}.\end{split}$ (48)
By (48), the function $g$ in (46) has a minimum at $\frac{K}{\mu^{2}}$ and
$g\left(\frac{K}{\mu}+1\right)=\log\left(\frac{K}{\mu}+1\right)-\frac{K}{\mu}<0,$
for all $K,\mu>0$. By (47) and (48), there exists $x>\frac{K}{\mu^{2}}$ such that (46) is strictly positive. Therefore, there exists $t^{\star}>\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)$ such that (45) holds. ∎
Appendix C Proof of the Main Results
We provide an upper bound for the moments of $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ defined in (8) and an estimate for the one step error associated with $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ which will be used in the proofs of Theorem 3.12 and Theorem 3.14 below.
Lemma C.1.
[7, Lemma 20] Let Assumptions 1 and 3.a hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
$\sup_{0\leq s\leq t}\mathbb{E}\left[|\widehat{Y}_{s}^{\text{EM}}|^{p}\right]\leq C_{\mathsf{EM},p}(t),$
where
$\begin{split}C_{\mathsf{EM},p}(t)&:=e^{t(3p-1-\frac{2}{p}+2^{2p-1}\mathsf{K}^{p}_{\text{Total}}(1+T^{\alpha p}))}\\ &\quad\times\left(\mathbb{E}\left[|\widehat{Y}_{0}^{\text{EM}}|^{p}\right]+2^{3p-2}\mathsf{K}^{p}_{\text{Total}}t(1+\mathbb{E}[|\hat{\theta}|^{p}])(1+T^{\alpha p})+\frac{2}{p}(pd+p(p-2))^{\frac{p}{2}}t\right).\end{split}$
Lemma C.2.
[7, Lemma 21] Let Assumptions 1 and 3.a hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
$\mathbb{E}\left[|\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}_{\lfloor t/\gamma\rfloor\gamma}^{\text{EM}}|^{p}\right]\leq\gamma^{\frac{p}{2}}C_{\mathsf{EMose},p},$
where
$\begin{split}C_{\mathsf{EMose},p}&:=2^{p-1}(C_{\mathsf{EM},p}(T)+\mathsf{K}^{p}_{\text{Total}}(1+T^{\alpha p})(2^{3p-2}C_{\mathsf{EM},p}(T)+2^{4p-3}(1+\mathbb{E}[|\hat{\theta}|^{p}])))\\ &\quad+(dp(p-1))^{\frac{p}{2}},\end{split}$
and $C_{\mathsf{EM},p}$ is defined in Lemma C.1.
Proof of Theorem 3.12.
We derive the non-asymptotic estimate for $W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ using the splitting
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))+W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))\\ &\quad+W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))+W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}})).\end{split}$ (49)
We provide upper bounds on the error made by the early stopping, i.e. $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))$, the error made by approximating the initial condition of the backward process $Y_{0}\sim\mathscr{L}(X_{T})$ with $\widetilde{Y}_{0}\sim\pi_{\infty}$, i.e. $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))$, the error made by approximating the score function with $s$, i.e. $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))$, and the discretisation error, i.e. $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}}))$, separately.
Upper bound on $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}})).$
The proof of this bound is the same as in the same bound provided in the [7, Proof of Theorem 10] using the representation of the OU process
[…]
$\begin{split}W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))&\leq 2\sqrt{\epsilon}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d}),\end{split}$ (51)
where $t_{K}=T-\epsilon$.
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}})).$
Using Itô’s formula, we have, for any $t\in[0,T-\epsilon]$,
$\begin{split}\text{d}|Y_{t}-\widetilde{Y}_{t}|^{2}&=2\langle Y_{t}-\widetilde{Y}_{t},Y_{t}+2\nabla\log p_{T-t}(Y_{t})-\widetilde{Y}_{t}-2\nabla\log p_{T-t}(\widetilde{Y}_{t})\rangle\ \text{d}t\\ &=2|Y_{t}-\widetilde{Y}_{t}|^{2}\ \text{d}t+4\langle Y_{t}-\widetilde{Y}_{t},\nabla\log p_{T-t}(Y_{t})-\nabla\log p_{T-t}(\widetilde{Y}_{t})\rangle\ \text{d}t.\end{split}$ (52)
By integrating and taking on both sides in (52), we have
$\begin{split}\mathbb{E}\left[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}\right]&=\mathbb{E}\left[|Y_{0}-\widetilde{Y}_{0}|^{2}\right]+\int_{0}^{t_{K}}2\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\ \mathrm{d}t\\ &\quad+\int_{0}^{t_{K}}4\mathbb{E}\left[\langle Y_{t}-\widetilde{Y}_{t},\nabla\log p_{T-t}(Y_{t})-\nabla\log p_{T-t}(\widetilde{Y}_{t})\rangle\right]\ \text{d}t.\end{split}$ (53)
[…]
$\begin{split}&\mathbb{E}\left[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}\right]\\ &\leq\mathbb{E}\left[|Y_{0}-\widetilde{Y}_{0}|^{2}\right]+2\int_{0}^{t_{K}}\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\ \mathrm{d}t-4\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\text{d}t\\ &\leq\mathbb{E}[|Y_{0}-\widetilde{Y}_{0}|^{2}]e^{2[t_{K}-2\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]}\\ &=\mathbb{E}[|m_{T}X_{0}+(\sigma_{T}-1)\widetilde{Y}_{0}|^{2}]e^{2[t_{K}-2\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]}\\ &\leq 2\left(\mathbb{E}[|X_{0}|^{2}]+d\right)\ e^{2[t_{K}-2\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]-2T}.\end{split}$ (54)
Using (54), Remark 3.9, and and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}]}\\ &\leq\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}.\end{split}$ (55)
Upper bound on $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}})).$
Using Itô’s formula, we have, for $t\in[0,T-\epsilon]$,
$\begin{split}\text{d}|\widetilde{Y}_{t}-Y_{t}^{\text{aux}}|^{2}&=2\langle\widetilde{Y}_{t}-Y^{\text{aux}}_{t},\widetilde{Y}_{t}+2\ \nabla\log p_{T-t}(\widetilde{Y}_{t})-Y^{\text{aux}}_{t}-2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}})\rangle\ \text{d}t\\ &=2|\widetilde{Y}_{t}-Y^{\text{aux}}_{t}|^{2}\ \text{d}t+4\ \langle\widetilde{Y}_{t}-Y^{\text{aux}}_{t},\nabla\log p_{T-t}(\widetilde{Y}_{t})-\nabla\log p_{T-t}(Y_{t}^{\text{aux}})\rangle\ \text{d}t\\ &\quad+4\ \langle\widetilde{Y}_{t}-Y^{\text{aux}}_{t},\nabla\log p_{T-t}(Y_{t}^{\text{aux}})-s(T-t,\hat{\theta},Y_{t}^{\text{aux}})\rangle\ \text{d}t.\end{split}$ (56)
By integrating and taking the expectation on both sides in (56), using Corollary 3.8, Young’s inequality with $\zeta\in(0,1)$ and Assumption 4, we have
$\begin{split}\mathbb{E}[|\widetilde{Y}_{T-\epsilon}-Y_{T-\epsilon}^{\text{aux}}|^{2}]&=2\int_{0}^{T-\epsilon}\mathbb{E}[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}]\ \text{d}s\\ &\quad+4\int_{0}^{T-\epsilon}\mathbb{E}[\langle\widetilde{Y}_{s}-Y^{\text{aux}}_{s},\nabla\log p_{T-s}(\widetilde{Y}_{s})-\nabla\log p_{T-s}(Y_{s}^{\text{aux}})\rangle]\ \text{d}s\\ &\quad+4\int_{0}^{T-\epsilon}\mathbb{E}[\langle\widetilde{Y}_{s}-Y^{\text{aux}}_{s},\nabla\log p_{T-s}(Y_{s}^{\text{aux}})-s(T-s,\hat{\theta},Y_{s}^{\text{aux}})\rangle]\ \text{d}s\\ &\leq\int_{0}^{T-\epsilon}2(1+\zeta)\ \mathbb{E}[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}]\ \text{d}s\\ &\quad-4\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-s}\mathbb{E}\left[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}\right]\text{d}t+2\zeta^{-1}\varepsilon_{\text{SN}}\\ &\leq 2e^{2(1+\zeta)(T-\epsilon)-4\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t}\ \zeta^{-1}\varepsilon_{\text{SN}}.\end{split}$ (57)
Using (57), Remark 3.9, and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y^{\text{aux}}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{K}}-Y_{t_{K}}^{\text{aux}}|^{2}]}\\ &\leq\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}.\end{split}$ (58)
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(\widehat{Y}_{t}^{\text{EM}})).$
Using Itô’s formula, we have, for $t\in[0,T-\epsilon]$,
$\begin{split}&\text{d}|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}\\ &=2\langle Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}},Y_{t}^{\text{aux}}+2\ s(T-t,\hat{\theta},Y_{t}^{\text{aux}})-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}-2\ s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma})\rangle\ \text{d}t\\ &=2|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}\ \text{d}t+2\langle Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}},\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}\rangle\ \text{d}t\\ &\quad+4\langle Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}},s(T-t,\hat{\theta},Y_{t}^{\text{aux}})-s(T-t,\hat{\theta},\widehat{Y}^{\text{EM}}_{t})\rangle\ \text{d}t\\ &\quad+4\langle Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}},s(T-t,\hat{\theta},\widehat{Y}^{\text{EM}}_{t})-s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma})\rangle\ \text{d}t.\end{split}$ (59)
Integrating and taking the expectation on both sides in (59), using Young’s inequality for $\zeta\in(0,1)$, Cauchy Schwarz inequality, Assumption 3.a, Lemma C.2, and Remark 3.1, we have
$\begin{split}\mathbb{E}\left[|Y_{T-\epsilon}^{\text{aux}}-\widehat{Y}_{T-\epsilon}^{\text{EM}}|^{2}\right]&\leq(2+3\zeta)\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t+\zeta^{-1}\int_{0}^{T-\epsilon}\mathbb{E}[|\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\ \text{d}t\\ &\quad+4\mathsf{K}_{3}(1+2T^{\alpha})\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\mathrm{d}t\\ &\quad+2\zeta^{-1}\int_{0}^{T-\epsilon}\mathbb{E}[|s(T-t,\hat{\theta},\widehat{Y}^{\text{EM}}_{t})-s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\mathrm{d}t\end{split}$
$\begin{split}&\leq(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t\\ &\quad+\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+4\mathbb{E}[|\hat{\theta}|^{2}])\\ &\quad+4\zeta^{-1}\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2}\int_{0}^{T-\epsilon}\mathbb{E}[|\widehat{Y}^{\text{EM}}_{t}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\mathrm{d}t\\ &\leq(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t\\ &\quad+\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}(1+4\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2})\\ &\quad+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})\\ &\leq e^{(2+3\zeta+4\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times\Bigg{(}\zeta^{-1}\gamma(T-\epsilon)C_{\mathsf{EMose},2}(1+4\mathsf{K}^{2}_{3}(1+2T^{\alpha})^{2})\\ &\qquad\quad+8\zeta^{-1}\gamma^{2\alpha}(T-\epsilon)\mathsf{K}_{1}^{2}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})\Bigg{)}.\end{split}$ (60)
Using (60) and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathscr{L}(Y_{T-\epsilon}^{\text{aux}}),\mathscr{L}(\widehat{Y}_{T-\epsilon}^{\text{EM}}))&\leq\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}).\end{split}$ (61)
Final upper bound on $W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting (51), (55), (58), and (61) into (49), we have
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})2\sqrt{\epsilon}\\ &\quad+\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}\\ &\quad+\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}\\ &\quad+\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}).\end{split}$ (62)
The bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ in (62) can be made arbitrarily small by appropriately choosing parameters including $\epsilon,T,\varepsilon_{\text{SN}}$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $0<\epsilon<\epsilon_{\delta}$ with $\epsilon_{\delta}$ given in Table 2 such that the first term on the right-hand side of (62) is
$(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})2\sqrt{\epsilon}<\delta/4.$ (63)
Next, we choose $T>T_{\delta}$ with $T_{\delta}$ given in Table 2 such that the second term on the right-hand side of (62) is
$\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}<\delta/4.$ (64)
Next, we turn to the third term on the right-hand side of (62). We choose $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ with $\varepsilon_{\text{SN},\delta}$ given in Table 2 such that
$\begin{split}\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}<\delta/4.\end{split}$ (65)
Finally, we choose $0<\gamma<\gamma_{\delta}$ with $\gamma_{\delta}$ given in Table 2 such that the fourth term on the right-hand side of (62) is
$\begin{split}&\gamma^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})<\delta/4.\end{split}$ (66)
Using (63), (64), (65), and (66), we obtain $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
Proof of Theorem 3.14.
Using the splitting (49), the proof follows along the same lines of the Proof of Theorem 3.12 for the estimation of the error bounds of the terms $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))$, $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))$, and $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))$. The error bound for $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}}))$ is derived in [7, Proof of Theorem 10]. Putting these four estimates together leads to (25) and (26). ∎
Appendix D Modified Half-Normal Distribution
In this section, we recall the probability density function of the modified half-normal distribution, see e.g. [59], used in Section 3.4.1 and defined as
$\begin{split}g(x)=\frac{2\xi^{\frac{\upsilon}{2}}x^{\upsilon-1}\exp\left(-\xi x^{2}+\psi x\right)}{\Psi\left(\frac{\upsilon}{2},\frac{\psi}{\sqrt{\xi}}\right)},\qquad x\geq 0,\end{split}$ (67)
where $\upsilon,\xi>0$, $\psi\in\mathbb{R}$, and the normalizing constant
$\begin{split}\Psi\left(\frac{\upsilon}{2},\frac{\psi}{\sqrt{\beta}}\right)&:=\sum_{n=0}^{\infty}\frac{\Gamma\left(\frac{\upsilon}{2}+\frac{n}{2}\right)}{\Gamma(n)}\frac{\psi^{n}\xi^{-n/2}}{n!},\end{split}$
is the Fox–Wright function [19, 66]. We point out that the half-normal distribution, truncated normal distribution, gamma distribution, and square root of the gamma distribution are all special cases of the modified Half-Normal distribution (67). The distribution (27) follows by taking the symmetric extension of (67), i.e. $g(|x|)/2$, and choosing $\upsilon=1$ and $\psi=-1$.
Appendix E Tables of Constants
Table 2 displays full expressions for constants which appear in Theorem 3.12 and Theorem 3.14.
Table 2: Explicit expressions for the constants in Theorem 3.12 and Theorem 3.14.
Constant Dependency Full Expression
$C_{1}$ $O(\sqrt{d})$ $2(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})$
$C_{2}$ $O(\sqrt{d})$ $\sqrt{2}\left(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d}\right)$
$C_{3}(T,\epsilon)$ $O(e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t})$ $\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}$
$C_{\mathsf{EM},2}(T)$ $O(Me^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $&e^{T(4+8\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))}\\ &\quad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{2}]+16\mathsf{K}^{2}_{\text{Total}}T(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})(1+T^{2\alpha})+2dT)\end{aligned}$
$C_{\mathsf{EM},4}(T)$ $O(d^{2}e^{T^{4\alpha+1}}T^{4\alpha+1})$ $&e^{T(\frac{21}{2}+128\mathsf{K}^{4}_{\text{Total}}(1+T^{4\alpha}))}\\ &\quad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{4}]+1024\mathsf{K}^{4}_{\text{Total}}T(1+\mathbb{E}[|\hat{\theta}|^{4}])(1+T^{4\alpha})+8(d^{2}+4d+4)T)\end{aligned}$
$C_{\mathsf{EMose},2}$ $O(de^{T^{2\alpha+1}}T^{4\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $2(C_{\mathsf{EM},2}(T)+\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(16C_{\mathsf{EM},2}(T)+32(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})))+2d\end{aligned}$
$C_{4}(T,\epsilon)$ $O(\sqrt{d}e^{T^{2\alpha+1}}T^{3\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})$ $&\zeta^{-1/2}(T-\epsilon)^{1/2}e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\qquad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})\end{aligned}$
$C_{\mathsf{EMose},4}$ $O(d^{2}e^{T^{4\alpha+1}}T^{8\alpha+1})$ $8(C_{\mathsf{EM},4}(T)+\mathsf{K}^{4}_{\text{Total}}(1+T^{4\alpha})(1024C_{\mathsf{EM},4}(T)+8192(1+\mathbb{E}[|\hat{\theta}|^{4}])))+144d^{2}\end{aligned}$
$\widetilde{C}_{4}(T,\epsilon)$ $O(de^{T^{4\alpha+1}}T^{4\alpha+1}\widetilde{\varepsilon}^{1/4}_{\text{AL}})$ $&\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\sqrt{T-\epsilon}\\ &\quad\ \times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\quad+4\zeta^{-1}d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)\\ &\qquad\qquad\qquad+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\times[d\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{1/2}\end{aligned}$
$\epsilon_{\delta}$ $\delta^{2}/(64(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})^{2})$
$T_{\delta}$ Obtained solving $T>T_{\delta}$ using Proposition 3.10, i.e., $&\ln(\mu(e^{2T}-1)+1)+(K/\mu+1)/(\mu(e^{2T}-1)+1)\\ &>\ln(4\sqrt{2}((\mathbb{E}[|X_{0}|^{2}])^{1/2}+\sqrt{d})/\delta)+2\int_{0}^{\epsilon}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t+K/\mu+1-\epsilon\end{aligned}$
$\varepsilon_{\text{SN},\delta}$ $(\delta^{2}\zeta/32)e^{-2(1+\zeta)(T-\epsilon)+4\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\end{aligned}$
$\gamma_{\delta}$ $&(\delta^{2}\zeta/16)(T-\epsilon)^{-1}e^{-2(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\quad\times(C^{1/2}_{\mathsf{EMose},2}(1+2\mathsf{K}_{3}(1+2T^{\alpha}))+2\sqrt{2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2})^{-2}\end{aligned}$
$\widetilde{\gamma}_{\delta}$ $\min\Bigg{\{}&(\delta/(4\sqrt{2}))^{1/\alpha}(T-\epsilon)^{-1/(2\alpha)}e^{-(2/\alpha)(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\\ &\times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad+4\zeta^{-1}d(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)\\ &\qquad\qquad+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\times[d\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{-1/(2\alpha)},1\Bigg{\}}\end{aligned}$